# On the original Ulam’s problem and its quantization

### v1

@@ Abstract.
We show that under general resonance the classical piecewise linear Fermi-Ulam accelerator behaves substantially different from its quantization, in the sense that the classical accelerator exhibits typical recurrence and non-escaping while the quantum version enjoys quadratic energy growth in general. We also describe a procedure to locate the escaping orbits, though exceptionally rare in the infinite-volume phase space, for the classical accelerators, which in particular include Ulam’s very original proposal and the linearly escaping orbits therein in the existing literature, and hence provide a complete (modulo a null set) answer to Ulam’s original question. For the quantum accelerators, we reveal under resonance the direct and explicit connection between the energy growth and the shape of the quasi-energy spectra.
Chern Institute of Mathematics and LPMC, Nankai University, Tianjin, China. Email: dongchg@nankai.edu.cn. School of Sciences, Great Bay University, Dongguan, China. Email: zhoujing@gbu.edu.cn
@@ 1. Introduction
In 1949 Fermi [Fermi49] proposed an acceleration mechanism, as to explain for the existence of high energy particles in cosmic rays, that charged particles get repeatedly reflected by moving magnetic mirrors, arguing that the chances of head-on interactions outnumber the head-tail ones so that the particles would on average get accelerated. Later in 1961 Ulam [Ulam61] extracted the effective mechanism in Fermi acceleration, i.e. a particle bounces between two infinitely heavy walls, one fixed and the other moving periodically (c.f. Fig.1). Ulam [Ulam61] performed numerical simulations with a piecewise linear wall motion and conjectured that such a model should produce escaping orbits, i.e. those whose energy grows to infinity in time.
@@
Since then, numerous efforts, theoretically and numerically, have been made by physicists and mathematicians in the study of the classical Fermi-Ulam accelerators and we refer to [Dol-FA, GRKT12, LiLi80] and references therein for detailed survey on the subject. The application of KAM theory leads to non-acceleration results for sufficiently smooth Fermi-Ulam accelerators and their variations [Dol08potential, kuOr20, LaLe91, Pu83, Pu94, Zhar2000]. On the other hand, escaping orbits appear if singularities are allowed [deSD12, Zha98, Zhou21], ambient potentials are introduced [AZ15, DS09, Pu77], and also higher-dimensions are considered [GT08nonauto, GT08ham, KZ23, STR10, Zhou20], where classical techniques such as Birkhoff normal forms and hyperbolic dynamics play fundamental roles.
@@
For the quantum Fermi-Ulam accelerators, the parallel non-acceleration results were obtained for generic sufficiently smooth wall motions [karn94, Seba90] with the help of quantum KAM theory [how89, how98]. Quantum chaos and acceleration behavior were expected both in numerics [glr14, joco86] and in theory [DoeRi69, Seba90] if singularities are allowed in the models. The quantum Fermi-Ulam accelerators are one of the earliest examples of quantum systems with time-dependent boundary conditions [DoeRi69], and have then inspired the development of many important techniques in the study of mesoscopic physics [Bas88, cbpwc96], quantum control [lbmw03, MaNa10], and nano-devices [AKM14, PiDBMS15], as well as experimental realizations such as cold atoms in dynamic traps [HalGMDHPN10].
@@
Notably among the existing literature on the subject, the closest related to our setting are the piecewise linear Fermi-Ulam accelerator [Zha98] and its quantization [Seba90]. In the piecewise linear classical Fermi-Ulam accelerator with the same particular choice of system parameters as Ulam [Ulam61], Zharnitsky [Zha98] found explicitly a family of linearly escaping orbits; however, these escaping orbits constitute a null set, leaving unanswered the behavior of the majority of orbits on the phase space of infinite volume. Besides, it was unclear therein [Zha98] if any other choices of parameters can also create acceleration. This issue of “mystery of lucky parameters” also appears in Šeba’s study of the piecewise linear quantum accelerators where he obtained quantum acceleration with the choice of parameters corresponding to the special $1:1$ resonance case in our definition (c.f. Definition LABEL:def:quantumresonanace and Example LABEL:example:11resonance).
@@
In this paper we study the class of piecewise linear Fermi-Ulam accelerators. We show that under general resonance conditions on the system parameters, the classical accelerators exhibit typical recurrence and the escaping orbits constitute a zero-measure set on the phase space of infinite volume. Meanwhile, the quantum resonant accelerators enjoy quadratic energy growth in general, and the quasi-energy spectra are absolutely continuous with finitely many continuity components. Our results cover and greatly generalize those in [Zha98] and [Seba90] as special cases.
@@
Also, we provide an explicit procedure on how to locate the escaping orbits, though exceptionally rare, in the classical resonant Fermi-Ulam accelerators. In particular, this includes Ulam’s very original proposal [Ulam61], and the linearly escaping orbits found by Zharnitsky therein [Zha98] correspond to an accelerating segment on a very particular invariant curve in our language, and hence we provide a complete answer (modulo a null set) to Ulam’s original question. In addition, for the quantum accelerators, we reveal the direct connection between the energy growth and the shape of the quasi-energy spectra with explicit formulas.
@@
From a technical point of view, in the classical accelerators, the rigidity of the piecewise linearity forces the dynamics to be singular and parabolic, making it impossible to apply classical tools such as KAM theory or hyperbolicity as done in many of the existing literature. For quantum accelerators, the time dependence, especially the delta jumps in wall motions, prevents direct application of a lot machinery for handling Schrödinger equations, just as in the case of the (in)famous kicked rotators. We combine several techniques from adiabatic theory, infinite ergodic theory, and quantum Floquet theory to tackle the challenges.
@@
The paper is organized as follows. We dissect the presentation of results on the classical piecewise linear Fermi-Ulam models in §2 and the corresponding quantum models in §LABEL:sec:QFU, and conclude with discussions on further problems in §LABEL:sec:conc. In §2, we begin with the adiabatic change of coordinates and the adiabatic normals, which serve as the Poincaré maps for the classical accelerators, and then we prove the typical recurrence results under resonance. In §LABEL:sec:QFU, we first transfer the time dependence in boundary conditions to time-dependent potentials in the Schrödinger equations, then we derive the Floquet wave propagators, and finally we prove the quadratic energy growth and the absolute continuity of the quasi-energy spectra under resonance.
@@ 2. The Classical Fermi-Ulam models
We consider the original Fermi-Ulam model, where a point particle bounces elastically between two infinitely heavy walls, one is fixed and the other moves periodically (c.f. Fig.1). Let $l(t)$ denote the distance between two walls. In Ulam’s original proposal [Ulam61], the periodic wall motion $l(t)$ is piecewise linear:
@@
The goal of this section is to show in Theorem LABEL:thm:rec that under resonance (as in Definition 2.5) the classical accelerators exhibit typical recurrence and the escaping orbits constitute a zero-measure set on the phase space of infinite volume. We also provide an explicit procedure for how to locate these exceptionally rare escaping orbits in the classical resonant Fermi-Ulam accelerators. In particular, this includes Ulam’s very original proposal [Ulam61] as a special resonance case, and the linearly escaping orbits found by Zharnitsky therein [Zha98] correspond to an accelerating segment on a very particular invariant curve (c.f. Example LABEL:example:escbdd).
The vital change of coordinates for the special parameters as in [Zha98] cannot be easily generalized, in the authors’ attempt, to other choices of parameters satisfying general resonance. Instead, we begin, in §2.1, with the adiabatic change of coordinates (c.f. Lemma 2.1) and the adiabatic normals (c.f. Proposition 2.2), which describe the behavior of the particle in roughly one period. Then in §2.2 we reveal the hidden skew product structure (c.f. Proposition LABEL:prop:skewprod) on the invariant circles (c.f. Lemma 2.4) due to resonance. Finally in §LABEL:ssec:rec we prove Theorem LABEL:thm:rec of the typical recurrence results for resonant accelerators.
@@ 2.1. The Adiabatic Normal Forms
In this section we present the Poincaré maps (c.f. Proposition 2.2) describing the momentum change of the particle in roughly one period $\Delta t=2T$ for large energies.
@@ 2.1.1. The Adiabatic Coordinates
suggesting that $I:=lv$ is almost constant. Then we update the algorithm by replacing $v$ with $I=lv$ and look for the next-order adiabatic invariant. On the other hand, the new “time” $\theta$ needs to behave in the way
@@
which leads to another ODE (we aim for constant $C=1$)
@@
Ideally, the Euler scheme above would terminate in finite steps since we are dealing with (piecewise) linear wall motions $l(t)$. We propose the following formulas for the adiabatic coordinates and verify in Lemma 2.1 that they are indeed adiabatic.

### v2

@@ Abstract.
We show that under general resonance the classical piecewise linear Fermi-Ulam accelerator behaves substantially different from its quantization, in the sense that the classical accelerator exhibits typical recurrence and non-escaping while the quantum version enjoys quadratic energy growth in general. We also describe a procedure to locate the escaping orbits, though exceptionally rare in the infinite-volume phase space, for the classical accelerators, which in particular include Ulam’s very original proposal and the linearly escaping orbits therein in the existing literature, and hence provide a complete (modulo a null set) answer to Ulam’s original question. For the quantum accelerators, we reveal under resonance the direct and explicit connection between the energy growth and the shape of the quasi-energy spectrum.
Chern Institute of Mathematics and LPMC, Nankai University, Tianjin, China. Email: dongchg@nankai.edu.cn. School of Sciences, Great Bay University, Dongguan, China. Email: zhoujing@gbu.edu.cn.
MSC classes: 37N05, 81Q50 (Primary), 35Q41, 37E10 (Secondary).
Keyword: Fermi acceleration, Fermi-Ulam model, resonance, escaping orbits, quasi-energy spectrum.
@@ 1. Introduction
In 1949 Fermi [Fermi49] proposed an acceleration mechanism, as to explain for the existence of high energy particles in cosmic rays, that charged particles get repeatedly reflected by moving magnetic mirrors, arguing that the chances of head-on interactions outnumber the head-tail ones so that the particles would on average get accelerated. Later in 1961 Ulam [Ulam61] extracted the effective mechanism in Fermi acceleration, i.e. a particle bounces between two infinitely heavy walls, one fixed and the other moving periodically (c.f. Fig.1). Ulam [Ulam61] performed numerical simulations with a specific piecewise linear wall motion (i.e. $A=1/\sqrt{2},B=\sqrt{2},T=1$ in Fig.1) and then conjectured that such a model should produce escaping orbits, i.e. those whose energy grows to infinity in time.
@@
Since then, numerous efforts, theoretically and numerically, have been made by physicists and mathematicians in the study of the classical Fermi-Ulam accelerators and we refer to [Dol-FA, GRKT12, LiLi80] and references therein for detailed survey on the subject. The application of KAM theory leads to non-acceleration results for sufficiently smooth Fermi-Ulam accelerators and their variations [Dol08potential, kuOr20, LaLe91, Pu83, Pu94, Zhar2000]. On the other hand, escaping orbits appear when singularities are allowed [deSD12, Zha98, Zhou21], ambient potentials are introduced [AZ15, DS09, Pu77], and also when higher-dimensions are considered [GT08nonauto, GT08ham, KZ23, STR10, Zhou20]. Classical techniques such as Birkhoff normal forms and hyperbolic dynamics play fundamental roles in the study of classical Fermi-Ulam accelerators.
@@
For the quantum Fermi-Ulam accelerators, the parallel non-acceleration results were obtained for generic sufficiently smooth wall motions [karn94, Seba90], with the help of quantum KAM theory [how89, how98]. Quantum chaos and acceleration behavior were expected both in numerics [glr14, joco86] and in theory [DoeRi69, Seba90] when singularities are allowed in the models. The quantum Fermi-Ulam accelerators are one of the earliest examples of quantum systems with time-dependent boundary conditions [DoeRi69], and have then inspired the development of many important techniques in the study of mesoscopic physics [Bas88, cbpwc96], quantum control [lbmw03, MaNa10], and nano-devices [AKM14, PiDBMS15], as well as experimental realizations such as cold atoms in dynamic traps [HalGMDHPN10].
@@
Notably among the existing literature on the subject, the closest related to our setting are a special case of piecewise linear Fermi-Ulam accelerator [Zha98] and its quantization [Seba90]. In a specific piecewise linear classical Fermi-Ulam accelerator with the same particular choice of system parameters as Ulam [Ulam61] (i.e. $A=1/\sqrt{2},B=\sqrt{2},T=1$ in Fig.1), Zharnitsky [Zha98] found explicitly a family of linearly escaping orbits. However, these escaping orbits constitute a null set, leaving unanswered the behavior of the majority of orbits on the phase space which has infinite volume. Besides, it was unclear therein [Zha98] if any other choices of parameters can also create acceleration. This issue of “mystery of lucky parameters” also appears in Šeba’s study of the piecewise linear quantum accelerators where he obtained quantum acceleration with the choice of parameters corresponding to the special $1:1$ resonance case in our definition (c.f. Definition LABEL:def:quantumresonanace and Example LABEL:example:11resonance).
@@
In this paper we study a general class of piecewise linear Fermi-Ulam accelerators. We show that under general resonance conditions on the system parameters, the classical accelerators exhibit typical recurrence and the escaping orbits constitute a zero-measure set on the phase space of infinite volume. Meanwhile, the quantum accelerators under general resonance enjoy quadratic energy growth in general, if not always, and the quasi-energy spectra are absolutely continuous with finitely many components. Our results cover and greatly generalize those in [Zha98] and [Seba90] as special cases.
@@
Also, we provide an explicit procedure on how to locate the escaping orbits, though exceptionally rare, in the classical Fermi-Ulam accelerators under general resonance. In particular, this includes Ulam’s very original proposal [Ulam61] (which is the special resonance case in our Definition 2.5), and the linearly escaping orbits found by Zharnitsky therein [Zha98] correspond to an accelerating segment on a very particular rational invariant curve in our language (c.f. Example LABEL:example:escbdd), and hence we provide a complete answer (modulo a null set) to Ulam’s original question. In addition, for the quantum accelerators, we reveal the direct connection between the energy growth behavior and the shape of the quasi-energy spectra with explicit formulas.
@@
From a technical perspective, in the classical accelerators, the rigidity of the piecewise linearity forces the dynamics to be singular and parabolic, making it impossible to apply classical tools such as KAM theory or hyperbolicity as has done in many of the existing literature. For quantum accelerators, the time dependence, especially the delta kicks in wall motions, prevents direct application of a lot machinery for handling Schrödinger equations, just as in the case of the (in)famous kicked rotators. We combine several techniques from adiabatic theory, infinite ergodic theory, and quantum Floquet theory to tackle the challenges.
@@
The paper is organized as follows. We dissect the presentation of results on the classical piecewise linear Fermi-Ulam models in §2 and the corresponding quantum models in §LABEL:sec:QFU, and conclude with discussions on open problems in §LABEL:sec:conc. In §2, we begin with the adiabatic change of coordinates and the adiabatic normals, which serve as the Poincaré maps for the classical accelerators, and then we prove the typical recurrence results under general resonance. In §LABEL:sec:QFU, we first transfer the time dependence in boundary conditions to time-dependent potentials in the Schrödinger equations, then we derive the Floquet wave propagators, and finally we prove the quadratic energy growth and the absolute continuity of the quasi-energy spectra under general resonance.
@@ 2. The Classical Fermi-Ulam models
We consider the original Fermi-Ulam model [Ulam61], where a point particle bounces elastically between two infinitely heavy walls, one is fixed and the other moves periodically (c.f. Fig.1). Let $l(t)$ denote the distance between two walls. In Ulam’s original proposal [Ulam61], the periodic wall motion $l(t)$ is piecewise linear:
@@
The goal of this section is to show in Theorem LABEL:thm:rec that under general resonance (as in Definition 2.5) the classical accelerators exhibit typical recurrence and the escaping orbits constitute a zero-measure set on the phase space of infinite volume. We also provide an explicit procedure for how to locate these exceptionally rare escaping orbits in the classical Fermi-Ulam accelerators under general resonance. In particular, this includes Ulam’s very original proposal (i.e. $A=1/\sqrt{2},B=\sqrt{2},T=1$) [Ulam61] as a particular case of the special resonance case, and the linearly escaping orbits found by Zharnitsky therein [Zha98] correspond to an accelerating segment on a very particular rational invariant curve (c.f. Example LABEL:example:escbdd).
The vital change of coordinates for the special parameters (i.e. $A=1/\sqrt{2},B=\sqrt{2},T=1$) as in [Zha98] cannot be easily generalized, in the authors’ attempt, to other choices of parameters satisfying general resonance. Instead, we begin, in §2.1, with the adiabatic change of coordinates (c.f. Lemma 2.1) and the adiabatic normals (c.f. Proposition 2.2), which describe the behavior of the particle in roughly one period. Then in §2.2 we show that the adiabatic normal forms possess invariant circles (c.f. Lemma 2.4) and then we reveal the hidden skew product structure (c.f. Proposition LABEL:prop:skewprod) on each invariant circle due to general resonance. Finally in §LABEL:ssec:rec we prove Theorem LABEL:thm:rec of the typical recurrence results for resonant accelerators and also describe an explicit procedure to locate escaping orbits on the rational invariant circles.
@@ 2.1. The Adiabatic Normal Forms
In this section the main goal is to derive the adiabatic normal forms (c.f. Proposition 2.2) which describe the momentum change of the particle in roughly one period $\Delta t=2T$ for large energies.
@@ 2.1.1. The Adiabatic Coordinates
suggesting that $I:=lv$ is almost constant. Then we update the algorithm by replacing $v$ with $I=lv$ and look for the next-order adiabatic invariant. On the other hand, the new “time” $\theta$ needs to behave in the following way
@@
which leads to another ODE (we aim for constant $C=1$ for simplicity)
@@
Ideally, the Euler scheme above would terminate in finite steps since we are dealing with (piecewise) linear wall motions $l(t)$. We propose the following formulas for the adiabatic coordinates and verify in Lemma 2.1 that this is indeed true.